# Generating a Paracosm for Training-Free Zero-Shot Composed Image Retrieval

### v1

@@ F Detailed Quantitative Results
We provide additional results in this section to further support the findings presented in the main paper. We include detailed numerical values corresponding to the radar plot (Fig.˜1) in Tab.˜7, offering a quantitative basis for comparing different methods across datasets. Secondly, we provide detailed results for each subset on GeneCIS in Tab.˜9, supporting the averaged results in Tab.˜3. In addition to the ablation study shown in Tab.˜5, we extend the analysis with two additional variants: using only synthetic counterparts without candidate images and adding detailed descriptions of candidate images in Tab.˜10. Tabs.˜11, 12 and 13 report benchmarking results on datasets using OpenCLIP ViT-G/14 [18] as the backbone. Under this unified setting, our method achieves the best or second-best performance on all datasets.
@@
Table 10: Ablation study. The results in this table are obtained on CLIP ViT-B/32. To evaluate the role of each component in our framework, we conduct a comprehensive ablation study by removing one module at a time while keeping the rest unchanged. based on candidate images. The red part in the table means images and texts from multimodal queries or generated based on multimodal queries. The blue part means images and texts from database or generated based on candidate images. $({\bf I}_{mental},{\mathbf{t}}_{query},{\mathbf{t}}_{mod})$ means (mental image, corresponding text description, modification text). $({\bf I}^{i},{\bf I}_{syn}^{i},{\mathbf{t}}^{i}_{detailed})$ means (candidate image, synthetic counterpart, detailed description of candidate image). Paracosm’s results are highlighted in bold.

### v2

@@ F Detailed Quantitative Results
We provide additional results in this section to further support the findings presented in the main paper. We include detailed numerical values corresponding to the radar plot (Fig.˜1) in Tab.˜8, offering a quantitative basis for comparing different methods across datasets. Secondly, we provide detailed results for each subset on GeneCIS in Tab.˜10, supporting the averaged results in Tab.˜3. In addition to the ablation study shown in Tab.˜6, we extend the analysis with two additional variants: using only synthetic counterparts without candidate images and adding detailed descriptions of candidate images in Tab.˜11. Tabs.˜12, 13 and 14 report benchmarking results on datasets using OpenCLIP ViT-G/14 [21] as the backbone. Under this unified setting, our method achieves the best or second-best performance on all datasets.
@@
Table 11: Ablation study. The results in this table are obtained on CLIP ViT-B/32. To evaluate the role of each component in our framework, we conduct a comprehensive ablation study by removing one module at a time while keeping the rest unchanged. The red part in the table means images and texts from multimodal queries or generated based on multimodal queries. The blue part means images and texts from database or generated based on candidate images. $({\bf I}_{mental},{\mathbf{t}}_{query},{\mathbf{t}}_{mod})$ means (mental image, corresponding text description, modification text). $({\bf I}^{i},{\bf I}_{syn}^{i},{\mathbf{t}}^{i}_{detailed})$ means (candidate image, synthetic counterpart, detailed description of candidate image). Paracosm’s results are highlighted in bold.